OpenStudy (anonymous):

Help! -6=2/13*x Solve the Equation im filling in my study guide for math and i cant get past this equation

OpenStudy (haseeb96):

Just apply the cross multiplication method here. That will be -6 will multiply with 13x=2

OpenStudy (haseeb96):

-72x=2 x=2/-72 x=-1/36 is the answer.

OpenStudy (anonymous):

thanks.. someone else sent me -39...

OpenStudy (freckles):

well i think it depends on what equation you have actually written

OpenStudy (freckles):

\[-6=\frac{2}{13}x ?\]

OpenStudy (freckles):

I think haseeb translated as \[-6=\frac{2}{13x}\] But I think -6(13) is actuallly -78 not -72

OpenStudy (anonymous):

the question you typed first is the exact one @freckles

OpenStudy (freckles):

do you know what the reciprocal of 2/13 is?

OpenStudy (anonymous):

by guessing id say 13/2 if i remembered the def right

OpenStudy (freckles):

then multiply both sides by 13/2